The following is the sales budget for Shleifer, Inc., for the first quarter of 2017:

January February March
Sales budget $ 216,000 $ 236,000 $ 259,000

  

Credit sales are collected as follows:
  
65 percent in the month of the sale.
30 percent in the month after the sale.
5 percent in the second month after the sale.
  
The accounts receivable balance at the end of the previous quarter was $100,000 ($70,000 of which was uncollected December sales).
  
a. Calculate the sales for November. (Do not round intermediate calculations and round your answer to the nearest whole dollar, e.g., 32.)
  
November sales           $  
  
b. Calculate the sales for December. (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
  
December sales           $  
  
c. Calculate the cash collections from sales for each month from January through March. (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)
   

Cash collections
January $
February $
March $
Answer #1

Solution-

a. November sales = (100000-70000)/5%= $6,00,000

b. December sales = 70000/(30%+5%) = $2,00,000

c. Cash collections :-

January = 600000*5%+200000*30%+216000*65%= $2,30,400

February = 200000*5%+216000*30%+ 236000*65%= $2,28,200

March = 216000*5%+236000*30%+259000*65%= $2,49,950